OpenStudy (anonymous):

how do I solve: ln(x)(3t+1)=ln(y)(2t+1) for y?

OpenStudy (campbell_st):

use change of base either base e or base 10 the change for base formula is \[\log_{x}(b) = \frac{\ln(b)}{\ln(x)}\] once you use change on base all terms are the same base so you then start to solve for y

OpenStudy (phi):

first step, divide both sides by (2t+1)

OpenStudy (phi):

then make each side the exponent of e

OpenStudy (john_es):

\[\ln(x^{3t+1})=\ln(y^{2t+1})\Rightarrow x^{3t+1}=y^{2t+1}\Rightarrow y=x^{\frac{3t+1}{2t+1}}\]

OpenStudy (anonymous):

i have \[\ln(x)*(1+3t) = \ln(y)*(1+2t)\] so you say I should do make that \[\frac{ \ln(x)(1+3t) }{ (1+3t)}=\ln(y)\]

OpenStudy (anonymous):

sorry (1+2t) on the bottom

OpenStudy (phi):

now make each side the exponent of e use the rule \[ e^{\ln(y)}= y \] on the left side you get \[ e^{\ln(x) A} \] where A is (1+3t)/(1+2t) the left side can be written as \[ \left(e^{\ln(x)} \right)^A \] now use the same rule as you did for the right side

OpenStudy (phi):

you should get the answer John_ES has already posted up above.